# High-Fidelity Data-Driven Dynamics Model for Reinforcement Learning-based Magnetic Control in HL-3 Tokamak

### v1

@@ Abstract
The drive to control tokamaks, a prominent technology in nuclear fusion, is essential due to its potential to provide a virtually unlimited source of clean energy. Reinforcement learning (RL) promises improved flexibility to manage the intricate and non-linear dynamics of the plasma encapsulated in a tokamak. However, RL typically requires substantial interaction with a simulator capable of accurately evolving the high-dimensional plasma state. Compared to first-principle-based simulators, whose intense computations lead to sluggish RL training, we devise an effective method to acquire a fully data-driven simulator, by mitigating the arising compounding error issue due to the underlying autoregressive nature. With high accuracy and appealing extrapolation capability, this high-fidelity dynamics model subsequently enables the rapid training of a qualified RL agent to directly generate engineering-reasonable magnetic coil commands, aiming at the desired long-term targets of plasma current and last closed flux surface. Together with a surrogate magnetic equilibrium reconstruction model EFITNN, the RL agent successfully maintains a $100$-ms, $1$ kHz trajectory control with accurate waveform tracking on the HL-3 tokamak. Furthermore, it also demonstrates the feasibility of zero-shot adaptation to changed triangularity targets, confirming the robustness of the developed data-driven dynamics model. Our work underscores the advantage of fully data-driven dynamics models in yielding RL-based trajectory control policies at a sufficiently fast pace, an anticipated engineering requirement in daily discharge practices for the upcoming ITER device.
Keywords: data-driven, dynamics model, reinforcement learning, magnetic control, HL-3 tokamak
@@ 1 Introduction
[…]
Fig. 1: The overall architecture of the plasma current and shape control in the HL-3 tokamak. a, The deployment of PID and RL control systems on HL-3. The historical interactions between the PID and the HL-3 tokamak produce the dataset for learning the data-driven dynamics model, which serves as an environment for RL training. RL controller is tasked with controlling the plasma current ($I_{p}$) and six shape parameters (minor radius $a$, elongation $\kappa$, upper triangularity $\delta_{u}$, lower triangularity $\delta_{l}$, radial position $R$ and vertical position $Z$ of plasma geometric center), with the output actions consisting of voltage commands for $17$ magnetic coils of the HL-3 tokamak. b, Schematic diagram of the localized model predict planning, wherein the dynamics model refrains from autoregressive extension and guides responsive actions only. c, Schematic diagram of long-term trajectory control. Different from using the physics-based simulator in [9], our work utilizes the dynamics model autoregressively to facilitate training RL-based controller. d, The proposed training diagram of the dynamics model, including scheduling sampling, multi-task adaptive weight adjustment, self-attention, and model ensembling.
@@
In this paper, we aim to provide a data-driven dynamics model, that exclusively learns from the discharge logs collected through the interaction between PID controllers and the tokamak device. The dynamics model is designed to efficiently support the fast training and practical deployment of a long-term, low-level RL-based agent, capable of directly generating voltage commands for up to $17$ magnetic coils. The system architecture is shown in Fig. 1. As shown in Fig. 1a, EFITNN [38], a surrogate model of magnetic equilibrium reconstruction code EFIT [39, 40], is leveraged to meaningfully transform the high-dimensional magnetic measurements to lower-dimensional plasma shape information while maintaining the physical interpretability and supporting real-time deployment. Subsequently, a Long Short-Term Memory Network (LSTM)-based [41] dynamics model is developed to track the evolution of plasma shape and current, based on accessible zero-dimensional plasma parameters and executed actuations. Benefiting from effectively leveraging several useful techniques including scheduling sampling [42, 43], adaptive weight adjustment [44, 45], self-attention [46] and model ensembling [47], the trained dynamics model demonstrates exceptional generalization capability on unseen plasma current regions during training and sensitively captures shape variations in typical scenarios, such as transitioning from limiter to (advanced) divertor configuration [48], and soft ramp-down [49] during discharges. Unlike localized short-term planning [5, 12, 11] shown in Fig. 1b, this data-driven dynamics model enables fully RL-based $100$-ms, $1$-kHz low-level trajectory control, achieving the desired Last Closed Flux Surface (LCFS) [50] via six shape and position parameters; while as in Fig. 1c, the training of $250,000$ iterations completes within $22$ minutes on the customer-grade Nvidia GeForce RTX 4090 GPU, fully satisfying the anticipation for RL-based magnetic control in ITER. It even achieves zero-shot triangularity adjustment, validating the effectiveness and reliability of the data-driven dynamics model.
@@ 2.1 Construction of Data-Driven Dynamics Model
Fig. 2: The demonstration case to predict plasma shape and position parameters for shot #6065 with and without the involvement of Scheduled Sampling (SS). Compared with actual discharge data (black solid line), a model trained with scheduled sampling (red dashed line) more effectively mitigates the compounding error than that without scheduled sampling (blue dashed line). The shaded area represents the results of ensembling $5$ independent models.
To train a $1$-kHz RL-based control policy, it is essential to build a high-fidelity data-driven dynamics model, which can autoregressively produce accurate predictions for up to $25$ variates. These predictive variates encompass the plasma current, $6$ plasma shape and position parameters, and $18$ channels of feedforward coil currents, which together correspond to a $44$-dimensional indispensable input (Table 2) and support the training of an engineering-reasonable policy for controlling $17$ magnetic coil voltages. In comparison to yielding $100$ ms-averaged evolution [12], forecasting single-variate tearability [5], or only tuning the total power and torque injected from the neutral beams [36], the significantly higher resolution and a larger number of coupling actuators implies to pose considerable challenges and requires a more comprehensive design.
@@
[…]
Fig. 3: Accuracy of the dynamics model. a, Ablations for key techniques in terms of MAE. SS: Scheduled Sampling, EN: Ensemble, SA: Self-Attention, AW: Adaptive Weight. Each group depicted here is derived by incorporating the corresponding module onto the basis of the preceding group. For instance, the notation “+ EN” signifies the integration of the Ensemble into the LSTM model that already incorporates scheduled sampling (“+ SS”). b, Goodness-of-fit of the dynamics model in terms of the coefficient of determination $R^{2}$.
@@ 2.2 Prediction Accuracy under Abrupt Shape variations
[…]
Fig. 4: The accuracy of dynamics model to predict abrupt plasma shape variations. a, Shot #6221, an H-mode discharge with ELMs and a plasma current of $300$ kA. b, Shot #6063, an advanced divertor configuration discharge with a plasma current of $500$ kA, with advanced divertor configuration undergoing a complete transformation between $1300$ ms and $1500$ ms. c, Shot #6232, an ITB discharge with a plasma current of $600$ kA. d, The LCFS variations correspond to the transition from limiter to divertor shape and soft ramp-down phases for shot #6221 ($300$ kA) and shot #6232 ($600$ kA), respectively.
In the previous section, we demonstrated the capability of the data-driven dynamics model in learning the high-dimensional evolution. In this section, we switch to investigate the prediction accuracy under abrupt plasma shape variations, which are commonly encountered due to the transition from limiter to divertor shape, the advanced divertor configuration transformation, or the initialization of soft ramp-down. Taking the example of shot #6221, an H-mode discharge with edge localized modes (ELMs), Fig. 4d manifests the significant impact of these activities on the plasma LCFS. Specifically, following the shift from the limiter to the divertor configuration, the plasma’s elongation $\kappa$, upper triangularity $\delta_{u}$, and lower triangularity $\delta_{l}$ surge abruptly. Fortunately, as shown in Fig. 4a, the dynamics model sensitively captures this variation. Similar robustness in predicting significant fluctuations in elongation and horizontal displacement can be observed in the initial segment of the ramp-down phase, where magnetic control still dominates despite the plasma’s strong interaction with the wall producing significant non-magnetic effects. On the other hand, Fig. 4b and Fig. 4c validate the accuracy of the dynamics model to predict abrupt shape variations in both advanced divertor configuration and Internal Transport Barrier (ITB) discharges [51, 52].
@@ 2.3 Extrapolation Capability on Unseen Plasma Current
Fig. 5: MAE averaging over 30 shots of 700 kA discharges. The numbers in the brackets correspond to the number of $700$ kA discharges included in the training dataset.
Fig. 6: Test results of shot #4042 under the dynamics model from zero-shot (blue dashed line), $2$-shot (green dashdot line), and full-shot training (red dotted line). The shaded area represents the results of the model ensemble, while actual discharge data is given in a black solid line.
Consistent with the staged research plan in ITER to gradually increase plasma current ($I_{p}$) [15], it is crucial to assess the extrapolation capability of a dynamics model trained on lower-$I_{p}$ datasets when applied to higher $I_{p}$ plasma discharge scenario. Therefore, we investigate the marginal contributions of incorporating various numbers of additional $700$ kA discharges to fine-tune a dynamics model initially trained on data with plasma currents ranging from $300$ to $600$ kA.
@@ 2.4 RL-based On-Device Control Results
[…]
Fig. 7: Target tracking control results of RL on both simulation (a) and actual devices (b). a, Simulating control results of plasma current $I_{p}$, shape and position parameters $a$, $\kappa$, $\delta_{u}$, $\delta_{l}$, $R$, and $Z$ are achieved, with the handover time set at the $1500$ ms. b, in shot #6691, the experimental data (red solid line) and the post-experiment replay of the dynamics model (blue dashdot line), as well as the target (black dashed line).
In line with the long-term trajectory magnetic control in [9], we aim to learn an RL-based policy to generate direct coil voltage commands for target plasma shape. This policy is trained through interactions with an autoregressively evolving dynamics model. During training, an actor-critic-based Proximal Policy Optimization (PPO) [53] is leveraged for learning to achieve the desired plasma current and shape, which is identical to the discharge result between $1500$ and $2000$ ms of shot #6226 of the HL-3 tokamak, an H-mode discharge with ELMs and a plasma current of $300$ kA. In PPO [53], the “actor” utilizes a policy to generate physics-reasonable magnetic coil voltage commands (see Methods) and modulate the plasma to the configuration of interest, while the “critic” evaluates the effectiveness of the current policy by estimating state values or advantages for potential policy refinement. Fig. 7a demonstrates the target tracking accuracy of the well-trained PPO actor within $1500$ $2000$ ms’ simulations. The control error of the plasma current is $2.55$ kA, or equivalently $0.85$% of the target for a $300$ kA discharge. The minor radius control error is $0.34$ cm, which corresponds to $0.52$% of the $65$ cm minor radius characteristic of the HL-3 tokamak device. Additionally, the control errors for the geometric center coordinates, both $R$ and $Z$, are constrained within a maximum of $1$ cm. The results indicate that the RL agent, supported by a data-driven dynamics model, can effectively explore and assimilate strategies that are advantageous for achieving discharge control objectives.
@@
Fig. 8: Error between EFITNN and offline EFIT.
2.4.1 RL-based $100$-ms Magnetic Control for Target Tracking
To thoroughly evaluate the flexibility of RL-based target tracking, we adopt a progressive approach. Initially, we assess the agent’s target tracking ability and ensure it aligns with the simulations. For this purpose, the experimental target remains identical to the reference shot #6226 for PPO training and the PPO actor responds to the variations in plasma shape and position parameters diagnosed by EFITNN. Despite the consistency in targets, noticeable differences arise during the PID control due to the inherent uncertainties in tokamak discharge and simultaneously coupled operations for HL-3 maintenance and testing.
In shot #6691, alongside the traditional PID control, we conduct an RL-based magnetic control from $1600$ ms to $1700$ ms. During this period, a $10$-ms transition period is established to ensure a stable switch between both control modes. Fig. 7b provides the corresponding control results.
Throughout the $100$-ms RL control, the minor radius $a$, elongation $\kappa$, and lower triangularity $\delta_{l}$ all progress with fast correction and accurate results as expected. However, due to the excessive horizontal displacement and acceleration on the handover time, radial position $R$ is not corrected fully, despite the adjustments being in the proper direction. Similar observations apply to the upper triangularity $\delta_{u}$, which deviates rapidly before $1600$ ms and is corrected progressively during the RL-based control period. Regarding plasma current $I_{p}$ and vertical position $Z$, the measured parameters fluctuate around the targets with moderate control errors. These outcomes manifest the robustness of the on-device RL-based control. Additionally, a post-experiment replay of the dynamics model (see Methods) closely aligns with the experimental results, confirming the effective and reliable support for RL decision-making.

### v2

# High-Fidelity Data-Driven Dynamics Model for Reinforcement Learning-based Control in HL-3 Tokamak

@@ Abstract
The success of reinforcement learning (RL)-based control in tokamaks, an emerging technique for controlled nuclear fusion with improved flexibility, typically requires substantial interaction with a simulator capable of accurately evolving the high-dimensional plasma state. Compared to first-principle-based simulators, whose intense computations lead to sluggish RL training, we devise an effective method to acquire a fully data-driven simulator, by mitigating the arising compounding error issue due to the underlying autoregressive nature. With high accuracy and appealing extrapolation capability, this high-fidelity dynamics model subsequently enables the rapid training of a qualified RL agent to directly generate engineering-reasonable actuator commands, aiming at the desired long-term targets of plasma configuration. Together with a surrogate model for Equilibrium Fitting code based on neural network, named EFITNN, the RL agent successfully maintains a $400$-ms, 11 kHz trajectory control with accurate waveform tracking of plasma current and last closed flux surface on the HL-3 tokamak. Furthermore, it also demonstrates the feasibility of zero-shot adaptation to changed triangularity targets, confirming the robustness of the developed data-driven dynamics model. Our work underscores the advantage of fully data-driven dynamics models in yielding RL-based trajectory control policies at a sufficiently fast pace, an anticipated engineering requirement in daily discharge practices for the upcoming ITER device.
Keywords: data-driven, dynamics model, digital twin, reinforcement learning-based control, HL-3 tokamak
@@ 1 Introduction
Fig. 1: The overall architecture of the plasma current and shape control in the HL-3 tokamak. a, Deployment schematic for plasma control on the HL-3 tokamak, illustrating two control paradigms. The Conventional Control System utilizes a Proportional-Integral-Derivative (PID) controller. It operates on the error between desired targets and a measured plasma state derived from the direct diagnostic measurements $\mathbf{m}$ via the M-Matrix, which calculates the $R$ and $Z$ positions for control, to ultimately generate commands that contribute to the final actuator voltages $\mathbf{a}$. The Reinforcement Learning (RL) Control System employs a trained policy controller. This controller receives its state $\mathbf{o}$ from the diagnostic measurements $\mathbf{m}$ processed by the EFITNN model. This state $\mathbf{o}$ is composed of the plasma current $I_{\mathrm{p}}$ as well as six shape and position parameters: the minor radius $a$, elongation $\kappa$, upper triangularity $\delta_{\mathrm{u}}$, lower triangularity $\delta_{\mathrm{l}}$, and the radial and vertical positions, $R$ and $Z$, of the plasma geometric center. The controller then takes this state $\mathbf{o}$ and the desired target state $\mathbf{d}$ as inputs to directly generate the action $\mathbf{a}$, which represents the voltage commands for the $17$ magnetic coils. Historical data from the conventional system, including direct diagnostic measurements $\mathbf{m_{t}}$, derived state observations $\mathbf{o_{t}}$, and the executed actuator commands $\mathbf{a_{t}}$, are collected to train the data-driven dynamics model. This model then serves as a high-fidelity simulated environment for the offline training of the RL controller. b, The proposed training diagram of the dynamics model, including scheduling sampling, multi-task adaptive weight adjustment, self-attention, and model ensembling.
@@
In this paper, we aim to provide a data-driven dynamics model, that exclusively learns from the discharge logs collected through the interaction between PID controllers and the tokamak device. The dynamics model is designed to efficiently support the fast training and practical deployment of a long-term, low-level RL-based agent, capable of directly generating actuator commands. The system architecture is shown in Fig. 1. As shown in Fig. 1a, EFITNN [38], a surrogate model of magnetic equilibrium reconstruction code EFIT [39, 40], is leveraged to meaningfully transform the high-dimensional magnetic measurements to lower-dimensional plasma shape information while maintaining the physical interpretability and supporting real-time deployment. Subsequently, a Long Short-Term Memory Network (LSTM)-based [41] dynamics model is developed to track the evolution of plasma shape and current, based on accessible zero-dimensional plasma parameters and executed actuations. Benefiting from effectively leveraging several useful techniques including scheduling sampling [42, 43], adaptive weight adjustment [44, 45], self-attention [46] and model ensembling [47], the trained dynamics model demonstrates exceptional generalization capability on unseen plasma current regions during training and sensitively captures shape variations in typical scenarios, such as transitioning from limiter to (advanced) divertor configuration [48], and soft ramp-down [49] during discharges. Unlike localized short-term planning [5, 12, 11] shown in Supplementary Figure 1a, this data-driven dynamics model enables fully RL-based $400$-ms, 11-kHz low-level trajectory control of up to $17$ magnetic coils, achieving the desired Last Closed Flux Surface (LCFS) [50] via six shape and position parameters; while as in Supplementary Figure 1b, the training of $250,000$ iterations completes within $22$ minutes on the customer-grade Nvidia GeForce RTX 4090 GPU, fully satisfying the anticipation for RL-based magnetic control in ITER. It even achieves zero-shot triangularity adjustment, validating the effectiveness and reliability of the data-driven dynamics model.
@@ 2.1 Construction of Data-Driven Dynamics Model
Fig. 2: The demonstration case to predict plasma shape and position parameters for shot #6065 with and without the involvement of Scheduled Sampling. Compared with actual discharge data (black solid line), a model trained with scheduled sampling (red dashed line) more effectively mitigates the compounding error than that without scheduled sampling (blue dashed line). The shaded area represents the results of ensembling 55 independent models.
To train a $1$-kHz RL-based control policy, it is essential to build a high-fidelity data-driven dynamics model, which can autoregressively produce accurate predictions for up to $25$ variates. These predictive variates encompass the plasma current, $6$ plasma shape and position parameters, and $18$ channels of feedforward coil currents, which together correspond to a $44$-dimensional indispensable input (Supplementary Table 2) and support the training of an engineering-reasonable policy for controlling $17$ magnetic coil voltages. In comparison to yielding $100$ ms-averaged evolution [12], forecasting single-variate tearability [5], or only tuning the total power and torque injected from the neutral beams [36], the significantly higher resolution and a larger number of coupling actuators implies to pose considerable challenges and requires a more comprehensive design.
@@
Fig. 3: Accuracy of the dynamics model. a, Ablations for key techniques in terms of MAE. $S$: Scheduled Sampling, EN: Ensemble, SA: Self-Attention, AW: Adaptive Weight. Each group depicted here is derived by incorporating the corresponding module onto the basis of the preceding group. For instance, the notation “+ EN” signifies the integration of the Ensemble into the LSTM model that already incorporates scheduled sampling (“+ SS”). b, Goodness-of-fit of the dynamics model in terms of the coefficient of determination $R^{2}$. The color map indicates the density of data points and the red dashed line represents the line of perfect agreement.
@@ 2.2 Prediction Accuracy under Abrupt Shape variations
Fig. 4: The accuracy of dynamics model to predict abrupt plasma shape variations. In panels a-c, the model’s prediction (red dashed line) is compared with experimental data (black solid line). The shaded regions highlight key operational phases: the light blue region marks the transition from the limiter to divertor configuration, the light orange region indicates the advanced divertor configuration transformation phase and the light green region denotes the plasma current ramp-down phase. a, Shot #6221, an H-mode discharge with ELMs and a plasma current of $300$ kA. b, Shot #6063, an advanced divertor configuration discharge with a plasma current of $500$ kA, with advanced divertor configuration undergoing a complete transformation between $1300$ ms and $1500$ ms. c, Shot #6232, an ITB discharge with a plasma current of $600$ kA. d, The Last Closed Flux Surface (LCFS) variations correspond to the transition from limiter to divertor shape and soft ramp-down phases for shot #6221 ($300$ kA) and shot #6232 ($600$ kA), respectively.
In the previous section, we demonstrated the capability of the data-driven dynamics model in learning the high-dimensional evolution. In this section, we switch to investigate the prediction accuracy under abrupt plasma shape variations, which are commonly encountered due to the transition from limiter to divertor shape, the advanced divertor configuration transformation, or the initialization of soft ramp-down. Taking the example of shot #6221, an H-mode discharge with edge localized modes (ELMs), Fig. 4d manifests the significant impact of these activities on the plasma LCFS. Specifically, following the shift from the limiter to the divertor configuration, the plasma’s elongation $\kappa$, upper triangularity $\delta_{\mathrm{u}}$, and lower triangularity $\delta_{\mathrm{l}}$ surge abruptly. Fortunately, as shown in Fig. 4a, the dynamics model sensitively captures this variation. Similar robustness in predicting significant fluctuations in elongation and horizontal displacement can be observed in the initial segment of the ramp-down phase, where magnetic control still dominates despite the plasma’s strong interaction with the wall producing significant non-magnetic effects. On the other hand, Fig. 4b and Fig. 4c validate the accuracy of the dynamics model to predict abrupt shape variations in both advanced divertor configuration and Internal Transport Barrier (ITB) discharges [51, 52].
@@ 2.3 Extrapolation Capability on Unseen Plasma Current
Fig. 5: Model error on 700 kA discharges. Mean Absolute Error (MAE) averaging over $30$ shots of $700$ kA discharges. The numbers in the brackets correspond to the number of $700$ kA discharges included in the training dataset.
Fig. 6: Few-shot learning performance on an unseen discharge. Test results of shot #4042 under the dynamics model from zero-shot (blue dashed line), 22-shot (green dashdot line), and full-shot training (red dotted line). The shaded area represents the results of the model ensemble, while actual discharge data is given in a black solid line.
Consistent with the staged research plan in ITER to gradually increase plasma current ($I_{\mathrm{p}}$) [15], it is crucial to assess the extrapolation capability of a dynamics model trained on lower-$I_{\mathrm{p}}$ datasets when applied to higher $I_{\mathrm{p}}$ plasma discharge scenario. Therefore, we investigate the marginal contributions of incorporating various numbers of additional $700$ kA discharges to fine-tune a dynamics model initially trained on data with plasma currents ranging from $300$ to $600$ kA.
@@ 2.4 RL-based On-Device Control Results
Fig. 7: Target tracking control results of RL on both simulation and actual devices. a, Simulated control results of plasma current $I_{\mathrm{p}}$, shape and position parameters $a$, $\kappa$, $\delta_{\mathrm{u}}$, $\delta_{\mathrm{l}}$, $R$, and $Z$. The RL control results (red solid line) track the target waveform from shot #6226 (black dashed line), with the handover time set at the $1500$ ms. b, in shot #12781, the experimental data (red solid line) and the post-experiment replay of the dynamics model (blue dashdot line), as well as the target (black dashed line).
Table 1: Plasma Shape Parameters during RL Control for Shot #12781
Time (ms) Elongation ($\kappa$) Upper Triangularity ($\delta_{\mathrm{u}}$) Lower Triangularity ($\delta_{\mathrm{l}}$)
1600 1.30 0.316 0.518
1630 1.40 0.347 0.612
1700 1.46 0.339 0.611
1800 1.42 0.269 0.639
1900 1.42 0.229 0.661
2000 1.38 0.182 0.620
In line with the long-term trajectory magnetic control in [9], we aim to learn an RL-based policy to generate direct coil voltage commands for target plasma shape. This policy is trained through interactions with an autoregressively evolving dynamics model. During training, an actor-critic-based Proximal Policy Optimization (PPO) [53] is leveraged for learning to achieve the desired plasma current and shape, which is identical to the discharge result between $1500$ and $2000$ ms of shot #6226 of the HL-3 tokamak, an H-mode discharge with ELMs and a plasma current of $300$ kA. In PPO [53], the “actor” utilizes a policy to generate physics-reasonable magnetic coil voltage commands (see Methods) and modulate the plasma to the configuration of interest, while the “critic” evaluates the effectiveness of the current policy by estimating state values or advantages for potential policy refinement. Fig. 7a demonstrates the target tracking accuracy of the well-trained PPO actor within $1500$ - $2000$ ms’ simulations. The control error of the plasma current is $2.55$ kA, or equivalently 0.850.85% of the target for a $300$ kA discharge. The control error for minor radius $a$ is $0.34$ cm, which corresponds to 0.520.52% of the $65$ cm minor radius characteristic of the HL-3 tokamak device. Additionally, the control errors for the geometric center coordinates, both $R$ and $Z$, are constrained within a maximum of $1$ cm. The results indicate that the RL agent, supported by a data-driven dynamics model, can effectively explore and assimilate strategies that are advantageous for achieving discharge control objectives.
@@
Fig. 8: Error between EFITNN and offline EFIT. Histograms of the prediction error between the EFITNN model and the offline EFIT code for key plasma parameters over all time slices in the test set. The Mean Absolute Error (MAE) for each parameter is shown in the title of its respective panel.
2.4.1 RL-based $400$-ms Magnetic Control for Target Tracking
To thoroughly evaluate the flexibility of RL-based target tracking, we adopt a progressive approach. Initially, we assess the agent’s target tracking ability and ensure it aligns with the simulations. For this purpose, the experimental target remains identical to the reference shot #6226 for PPO training and the PPO actor responds to the variations in plasma shape and position parameters diagnosed by EFITNN. Despite the consistency in targets, noticeable differences arise during the PID control due to the inherent uncertainties in tokamak discharge and simultaneously coupled operations for HL-3 maintenance and testing. A series of RL control experiments have been performed on HL-3.
Typically, in shot #12781, alongside the traditional PID control, we conduct an RL-based magnetic control from $1600$ ms to $2000$ ms with $30$ ms transition periods at the start and end. Fig. 7b provides the corresponding control results and Table 1 details the evolution of elongation ($\kappa$) and upper and lower triangularities ($\delta_{\mathrm{u}}$ and $\delta_{\mathrm{l}}$) during the RL control phase. During its operational phase, the RL controller demonstrates excellent tracking performance for the four plasma shape parameters ($a$, $\kappa$, $\delta_{\mathrm{u}}$, and $\delta_{\mathrm{l}}$). It rapidly adjusts these parameters to their target values and subsequently maintains their stability with high precision. On par with PID, the controller capably maintains the radial ($R$) and vertical ($Z$) positions within a stable range, with precision slightly lower than that of shape parameters. Additionally, the control handover induces a brief fluctuation in the plasma current ($I_{\mathrm{p}}$), which the RL controller promptly corrects, restoring it to the target level. The experimental results confirm that the RL controller possesses the capability for precise plasma shape regulation. In contrast to the conventional PID system, which is limited to controlling plasma positions ($R$ and $Z$) and current ($I_{\mathrm{p}}$), the RL controller successfully extends the closed-loop scope to the simultaneous regulation of four additional key shape parameters ($a$, $\kappa$, $\delta_{\mathrm{u}}$, and $\delta_{\mathrm{l}}$).